Neveready Flashlights, Inc. needs $410,000 to take a cash discount of 2/10, net 60. A banker will loan the money for 50 days at an interest cost of $6,600.

a.

What is the annual rate on the bank loan? (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

  Annual rate %
b.

How much would it cost (in percentage terms) if the firm did not take the cash discount, but paid the bill in 60 days instead of 10 days? (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

  Cost of not taking a cash discount %
c. Should the firm borrow the money to take the discount?
  • Yes

  • No

d.

If the banker requires a 13 percent compensating balance, how much must the firm borrow to end up with the $410,000? (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

  Amount to be borrowed $   
e-1.

What would be the effective interest rate in part d if the interest charge for 50 days were $8,600? (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

  Effective interest rate %
e-2.

Should the firm borrow with the 18 percent compensating balance? (The firm has no funds to count against the compensating balance requirement.)

  • No

  • Yes
Answer #1

ANSWER

(a) Total Bill = $ 410000, Discount = 2 % Loan Tenure = 50 days, Total Interest = $ 6600

Annual Interest Rate = (6600/410000) x (365/50) x 100 = 11.75 %

(b) Cash Discount =2 % of bill = 0.02 x 410000 = $ 8200

If the firm does not take the discount and instead pays in 60 days, it is equivalent to losing $ 8200 on a payment of $ 410000 over a period of 60 day.

Annual Cost = (8200 / 410000) x (365/60) x 100 = 12.167 %

(c) As the cost of borrowing is lesser than the cost of foregoing the cash discount, the firm should borrow to avail the cash discount.

(d) Compensating Balance = 13 % of Total Borrowing

Total Borrowing - Compensating Balance = $ 410000

0.87 x Total Borrowing = $ 410000

Total Borrowing = 410000 / 0.87 = $ 471264.37

NOTE: Please raise separate queries for solutions to the remaining sub-parts.

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************